OpenStudy (anonymous):

show how to prove this identity. cot^(2)θ=cos^(2)θ+((cotθ)(cosθ))^(2)

OpenStudy (anonymous):

take the right hand side \[cos^2θ+((cotθ)(cosθ))^2\]take out cos^(2)θ as a common term \[cos^2θ[1+(cotθ)^2]\]Since\[1+(cotθ)^2 = Cosec^2θ\]plug it in and simplify to get cot^(2)θ

OpenStudy (anonymous):

did you understand the method?

OpenStudy (anonymous):

i cant understand this part. pls explain. thanks! cos2θ[1+(cotθ)2]

OpenStudy (anonymous):

Since \[Cosec^2\theta= \frac{1}{Sin^2\theta}\]\[Cos^2\theta*Cosec^2\theta=\frac{Cos^2\theta}{Sin^2\theta} = Cot^2\theta\]

OpenStudy (anonymous):

i got it. thanks a lot akitav!